Define equilibrium constant and for the following reaction


2NH↓3(g) ⇌ 2N↓2(g) +3H↓2 (g)


i. Write the expression for Kc.


ii. Relate Kc to Kp, and


iii. Relate Kc to Kx

1
Expert's answer
2022-02-21T17:07:07-0500

Solution:

Balanced chemical equation:

2NH3(g) ⇌ N2(g) + 3H2(g)


The expression for Kc:

Relationship between Kc and Kp:

Kp = Kc × (RT)Δn

Δn = (moles of gaseous product) − (moles of gaseous reactant)

Δn = 4 − 2 = 2

Hence,

Kp = Kc × (RT)2


Relationship between Kc and Kx:

Kx = Kc × (P / RT)Δn

Hence,

Kx = Kc × (P / RT)2